1. Find the equilibrium wage rate. This is done by setting the quantity demanded equal to the quantity supplied and algebraically solving for W. That is, solve the following equation for W:

24,000 – 320W = -18,000 + 1250W
Answer #1

Answer:

Answer: Based on given information to find W we have to set Qd = Qs. So, 24,000 - 320W = -18,000 + 1250W => 24,000 + 18,000 =